# DRDT3: Diffusion-Refined Decision Test-Time Training Model

### v1

@@ IV-C2 Ablation on Diffusion
IV-C3 Ablation on Noise Approximator
We also verify the effectiveness of our gated MLP noise approximator by comparing DRDT3 with several variants: one where the gated MLP structure is replaced by a simple MLP, another where the adaLN is replaced with in-context conditioning [27], and a third variant that removes both structures. As shown in Table V, the removal of adaLN causes a slight performance degradation, while removing the gated MLP results in a significant reduction in performance.
IV-C4 Ablation on Loss Function
To verify the effectiveness of our proposed unified optimization objective, we compare our loss function with the vanilla diffusion loss function (w/o $\mathcal{L}_{dt3}$) and a unified loss function incorporating an L2-based DT3 loss term (w/ L2 $\mathcal{L}_{dt3}$). As shown in Table V, the variant w/o $\mathcal{L}_{dt3}$ shows worse performance than the variant with an extra L2-based DT3 constraint. Our DRDT3, utilizing an L1-based DT3 loss term instead, can further enhance performance beyond the L2 variant. These results demonstrate that incorporating DT3 loss into the final loss improves performance, with the L1 $\mathcal{L}_{dt3}$ offering additional gains over the L2 variant.
IV-C5 Sensitivity Analysis on Loss Coefficient
@@
TABLE V: Ablation studies on the noise approximator (left side) and unified loss function (right side).
@@
V Related Work
V-A Decision Transformer
Offline RL learns policies from historical trajectories pre-collected using other behavior policies. Different from classic offline RL [34, 12] where dynamics programing is employed for policy optimization, DT [14] formulates the policy learning as a sequence modeling problem by autoregressively generating trajectories with Generative Pre-trained Transformer (GPT)-like architectures [16]. Many variants of DT have been studied and shown competitive performance on some RL benchmarks. Online DT [33] equips DT with entropy regularizers, enabling online finetuning with efficient exploration. Prompting DT [39] and Hyper-DT [40] make DT easily transferable to novel tasks using trajectory prompt and adaptation module initialized with hyper-network, respectively. Hierarchical DT [41] frees DT from specifying RTGs while using sub-goal selection instead. Q-learning DT [10] and Elastic DT [17] improve the stitching ability of DT by introducing dynamic programing-derived value function and varying-length context, respectively. DT has also been applied to some real-world applications, such as robotics [41] and traffic signal control [42].
V-B Sequence Modeling
Transformer-based methods leveraging the attention mechanism have demonstrated remarkable performance across various sequence modeling domains [15, 16, 43, 27]. However, their quadratic time complexity during inference [38] poses challenges for long-sequence modeling. Recently, structured state space models (SSMs) [44, 38] have introduced a compelling architecture for long-sequence processing, distinguished by their linear computational complexity. Among SSM-based methods, MAMBA has proven effective and is widely applied in various research areas [45, 46]. MAMBA [38] introduces a selection mechanism to adapt SSM parameters based on input, along with parallel scanning, kernel fusion, and recomputation techniques for efficient computation. Other recent sequence modeling approaches, such as RWKV [47, 48], Gated Linear Attention (GLA) [49], and xLSTM [50], build on RNN architectures with enhanced expressiveness by utilizing matrix hidden states, unlike the vector hidden states of traditional RNNs. Additionally, TTT layers [25] further enhance hidden state expressiveness by updating them through self-supervised learning during both training and inference. TTT-Linear and TTT-MLP, whose hidden states are a linear layer and an MLP respectively, achieve linear time complexity with performance that is comparable to, or even surpasses, Transformers and MAMBA.
V-C Diffusion Models in Reinforcement Learning
Diffusion models [18] have been appealing for their strong capacity to generate high-dimension image or text data. In light of the expressive representation and strong multi-modal distribution modeling ability of the diffusion model, some researchers introduce the diffusion model to RL paradigms. Diffuser [21] first employs diffusion models as a planner for generating trajectories in model-based offline RL, which alleviates the severe compounding errors of conventional planners. Diffusion-QL [9] explores representing policy as a diffusion model and employing Q-value function guidance during training. It overcomes the over-conservatism of policies learned from conventional offline RL. Decision Diffuser [22] formulates sequential decision-making problems as conditional generative modeling and introduces constraints and skill conditions. In addition to employing diffusion models as planners or policies, SynthER [51] further exploits diffusion models as a data synthesizer to augment both offline and online training data with high diversity.
Denoising Diffusion Probabilistic Models (DDPM) [18], a well-known diffusion model, samples data through an iterative denoising process. Our work proposes a unified framework that aims to enhance a DT-style trajectory modeling method by refining the predictions iteratively within the expressive DDPM.
VI Conclusion and Discussion
@@
This work introduces the DRDT3 algorithm to enhance the stitching ability of the DT-based method. We propose a novel DT3 module with attention TTT blocks to process historical information and predict a better coarse action representation. The diffusion module then conditions this coarse representation and iteratively refines the action using a novel gated MLP noise approximator, improving performance on sub-optimal datasets. To achieve joint learning of the DT3 and diffusion modules, we introduce a unified optimization objective, consisting of both an action representation term and an action refinement term. Experiments on a variety of tasks with different dataset configurations from the D4RL benchmark demonstrate the effectiveness and superior performance of DRDT3 compared to conventional offline RL methods and DT-based approaches. Our DT3 trained along also shows better performance compared to vanilla DT. In future work, we plan to explore the potential of extending DRDT3 for online adaptation.

### v2

@@ 5.2 Performance Comparison
Table 6: Performance comparison of DT, DT3, and DRDT3 across Adroit and Humanoid tasks.
Method pen-h pen-c door-h door-c Avg. human-m human-me human-e Avg.
DT 72.4 40.2 9.2 3.7 31.4 23.7 52.6 53.4 43.2
[…]
5.3 Ablation on Sequence Model
We compare our DRDT3 framework with two variants: one where the sequence model is replaced by a GPT-2 model (Radford et al., 2019) and another where the TTT layer is replaced by an MAMBA layer (Gu & Dao, 2023). The GPT-2 model relies entirely on the attention mechanism, while MAMBA is a selective state space model based on the RNN structure, where parameters are dynamically dependent on the inputs. For the two variants w/GPT-2 and w/ Attention MAMBA, we use one block per model as our DRDT3. To ensure a fair comparison, we adjust the number of layers in both the GPT-2 and MAMBA variants to match the model size of our DRDT3. Specifically, the total number of learnable parameters for our DRDT3, the variant w/GPT-2, and the variant w/ Attention MAMBA are 235K, 218K, and 329K, respectively.
As shown in Figure 3, our DRDT3 with Attention TTT blocks outperforms the other variants across three Medium-Expert tasks. The variant with Attention MAMBA shows slight performance degradation and lacks the stability of DRDT3. The variant with GPT-2 performs the worst across all tasks, which might be attributed to the typical reliance on larger model sizes of GPT-2. When constrained to a similar size as the other two variants, GPT-2 experiences performance degradation. These results indicate that our proposed DT3 module with Attention TTT blocks enhances performance without increasing model size.
Figure 4: Learning curves of DT, DT3, and DRDT3 on Gym tasks. For clarity, each curve is plotted with a moving average (window size = 10) to highlight the trend.
5.4 Ablation on Diffusion
To test whether the diffusion module plays a significant role in action refinement and verify the effectiveness of our DT3 module, we train a DT3 model along with the same settings as DRDT3 by minimizing the $\mathcal{L}_{dt3}$ in Equation 16. Table 4, 5, and 6 show the performance comparison between DT and our DT3 model on Gym, AntMaze, and challenging robotic (Adroit and Humanoid) tasks. Figure 4 illustrates the learning curves of the three methods on Gym locomotion tasks.
As we can see, our proposed DT3 trained independently can outperform DT on 6 out of 9 Gym locomotion tasks, all AntMaze tasks, 3 out of 4 Adroit tasks, and 2 out of 3 Humanoid tasks. This demonstrates the superiority of our proposed DT3 with the Attention TTT blocks over the vanilla DT with only the attention mechanism. Furthermore, our proposed DRDT3, which enhances DT3 through a diffusion module, further boosts performance across all tasks with different datasets, achieving average improvements of 10.3%, 7.6%, 38.0%, 27.7%, and 24.1% on Gym locomotion, Ant, AntMaze, Aroit, and Humanoid tasks, respectively. This demonstrates the substantial impact of incorporating diffusion refinement in DRDT3.
However, underperformance is observed in the walker2d-me task. As shown in Figure 4, we can see a slight degradation in the learning curve as training iterations increase, indicative of overfitting to the dataset’s mixed-quality dynamics. Specifically, as training progresses, the DT3 module might memorize noisy or erratic details from medium-quality trajectories rather than generalizing to the smooth, energy-efficient patterns of expert demonstrations. This leads to increasingly misaligned coarse actions, which the diffusion module cannot fully correct. These may result in degraded policy performance on this task.
Table 7: Model size and average training and inference time comparison. # param. denotes the number of learnable parameters. Training time refers to the average time required for one epoch of model training, while inference time indicates the time needed to evaluate the policy for a single episode.
Model Size Hopper-Medium HalfCheetah-Medium Walker2d-Medium
# param. (K) Training (s) Inference (s) Training (s) Inference (s) Training (s) Inference (s)
DT 226.8 162.2 1.7 258.7 2.6 239.1 2.6
DT3 234.6 134.3 2.6 245.8 4.2 223.9 3.1
5.5 Computational Efficiency
We evaluate the computational efficiency of our proposed DT3, which consists of Attention–TTT blocks, by comparing it against DT, which composed solely of Attention blocks. Table 7 reports the model sizes along with the average training and inference times. As shown, DT3 consistently achieves faster training but fails to surpass DT in inference speed. This is consistent with the findings in the original TTT paper (Sun et al., 2024), which reported that TTT layers offer a clear advantage over Transformers during training and become more efficient at inference when processing long contexts. In our case, the sequence length is only 6, a relatively short context, which prevents DT3 from benefiting from TTT’s inference-time speed advantage.
@@ 5.5 Computational Efficiency
Table 8: Ablation studies on the noise approximator (left side) and unified loss function (right side).
@@
5.6 Ablation on Noise Approximator
We also verify the effectiveness of our gated MLP noise approximator by comparing DRDT3 with several variants, as shown in Figure 5: one where the gated MLP structure is replaced by a simple MLP, another where the adaLN is replaced with in-context conditioning (Peebles & Xie, 2023), and a third variant that removes both structures. As shown in Table 8, the removal of adaLN causes a slight performance degradation, while removing the gated MLP results in a significant reduction in performance.
5.7 Ablation on Loss Function
To verify the effectiveness of our proposed unified optimization objective, we compare our loss function with the vanilla diffusion loss function (w/o $\mathcal{L}_{dt3}$) and a unified loss function incorporating an L2-based DT3 loss term (w/ L2 $\mathcal{L}_{dt3}$). As shown in Table 8, the variant w/o $\mathcal{L}_{dt3}$ shows worse performance than the variant with an extra L2-based DT3 constraint. Our DRDT3, utilizing an L1-based DT3 loss term instead, can further enhance performance beyond the L2 variant. These results demonstrate that incorporating DT3 loss into the final loss improves performance, with the L1 $\mathcal{L}_{dt3}$ offering additional gains over the L2 variant.
5.8 Sensitivity Analysis on Loss Coefficient
6 Conclusion and Discussion
@@
This work introduces the DRDT3 algorithm to enhance the stitching ability of the DT-based method. We propose a novel DT3 module with attention TTT blocks to process historical information and predict a better coarse action representation. The diffusion module then conditions this coarse representation and iteratively refines the action using a novel gated MLP noise approximator, improving performance on sub-optimal datasets. To achieve joint learning of the DT3 and diffusion modules, we introduce a unified optimization objective, consisting of both an action representation term and an action refinement term. Experiments on a variety of tasks with different dataset configurations from the D4RL benchmark demonstrate the effectiveness and superior performance of DRDT3 compared to conventional offline RL methods and DT-based approaches. Experiment results also demonstrate that DT3 can outperform the vanilla DT.
Nonetheless, we acknowledge that there remains a performance gap between our DRDT3 and certain state-of-the-art dynamic programming–assisted diffusion-based policies. This gap likely arises because DRDT3 is inherently constrained by the performance of the underlying DT3 module, as it directly refines the actions generated by DT3. Therefore, we plan to further investigate strategies to narrow this gap in the future, aiming to bring DT-based methods closer to the performance of their dynamic programming–based counterparts.